# APTx Neuron: A Unified Trainable Neuron Architecture Integrating Activation and Computation

### v1

@@ 3.4 Parameter Overhead and Efficiency
While the present formulation employs distinct parameters $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ for each input dimension $x_{i}$ within a neuron, a parameter-sharing scheme, where $\alpha_{i}=\alpha$, $\beta_{i}=\beta$, and $\gamma_{i}=\gamma$ across all input dimensions fed to the neuron, may be adopted to reduce model complexity. Such a variant may offer a trade-off between parameter efficiency and expressive power.
@@ 4.1 Neural Network Design
The APTx Neuron-based feedforward neural network replaces conventional linear and activation layers with custom layers composed of multiple APTx Neurons. Each APTx Neuron unifies computation and non-linearity into a single trainable expression. The architecture used in our experiments is as follows:

### v2

@@ 3.4 Parameter Overhead and Efficiency
Although the original formulation uses parameters $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ for each input dimension $x_{i}$ within a neuron, along with a shared parameter $\delta$, several parameter sharing strategies can be adopted to reduce model complexity:
  • Full sharing: All parameters are shared and trainable across input dimensions, i.e., $\alpha_{i}=\alpha$, $\beta_{i}=\beta$, $\gamma_{i}=\gamma$, and a shared, trainable $\delta$.
  • Partial sharing: For example, $\alpha_{i}=\alpha$ (shared and trainable), while $\beta_{i}$ and $\gamma_{i}$ remain input specific and trainable; $\delta$ remains shared and trainable.
  • Hybrid schemes: Certain parameters may be fixed. For instance, setting $\alpha_{i}=1$ (non-trainable), while keeping $\beta_{i}$, $\gamma_{i}$, and $\delta$ trainable. For a neuron with input dimension $n$, this reduces the total number of trainable parameters from $13n+$ 1 to $12n+$ 1.
These variants offer configurable trade-offs between parameter efficiency and expressive power, allowing flexible adaptation of the APTx Neuron to a variety of architectures and tasks.
@@ 4.1 Neural Network Design
The APTx Neuron-based feedforward neural network replaces conventional linear and activation layers with custom layers composed of multiple APTx Neurons. Each APTx Neuron unifies computation and non-linearity into a single trainable expression. In this configuration, all parameters $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, and $\delta$ were trainable. The architecture used in our experiments is as follows: